OpenStudy (ralph):

f(x)=(6x^2)-2x Find the limit. Deltax->0

OpenStudy (aabomosalam1998):

\[f \prime (x)=\lim_{ \Delta x \rightarrow 0} f(x+\Delta x)-f(x)/\Delta x=\lim_{\Delta x \rightarrow 0} 6(x+\Delta x)^2-2(x+\Delta x)-(6x^2-2x)/\Delta x\]

OpenStudy (raffle_snaffle):

Nice just give them the answer...

OpenStudy (ralph):

I've gotten that far but I don't understand how to factor it.

OpenStudy (ralph):

I understand you put it all over deltax but yeah

OpenStudy (ralph):

like that entire equation divided by deltax

zepdrix (zepdrix):

@AAbomosalam1998 try to get used to using the fraction command: \frac{}{} it will make your equations look a lot nicer :)

OpenStudy (aabomosalam1998):

Yes, that's right, after expanding the brackets and combining like terms, factor out delta x and then take the limit as delta x approaches zero. @zepdrix okay, I am sorry I am not professional with typing out the commands in latex, I will use that next time.

OpenStudy (ralph):

I don't know how to factor a problem with deltax in it, like squaring it, are there rules to it?

OpenStudy (aabomosalam1998):

6(x+Δx)^2=6(x+Δx)(x+Δx) Multiply each term in the first bracket with each term in the second bracket, add like terms then multiply each term by 6. Do the same to -2(x+Δx). 6x^2+12xΔ x-6Δx^2-2x-2Δx-6x^2+2x=12xΔx-6Δx^2-2Δx . Factor out Δx =12x-6Δ x-2. Take the limit as Δx-->0

OpenStudy (aabomosalam1998):

\[6x^2+12x \Delta x-6 \Delta x^2-2x-2\Delta x-6x^2+2x=12x \Delta x -6 \Delta x^2-2\Delta x\] \[ \frac{12x \Delta x-6 \Delta x ^2-2\Delta x}{\Delta x}=\frac{\Delta x(12x-6 \Delta x-2)}{\Delta x}=12x-6\Delta x-2\]

OpenStudy (ralph):

so the limit if x approaches 0 is -2?

OpenStudy (phi):

if the question is what is the limit as x ->0 of the expression 6x^2 -2x then the limit will be the sum of the limits of each term lim (6x^2) + lim(-2x) the limit of a product is the limit of the products so lim(6x^2) = lim(6) * lim(x^2) the limit of a constant is the constant: lim(6)*lim(x^2) = 6 * lim(x^2) any idea what \[ \lim_{x\rightarrow 0} x^2 \] ? it is 0 ditto for lim(-2x) = -2*lim(x)

OpenStudy (mathmale):

First of all, we need clarification regarding what you're trying to do here. f(x)=(6x^2)-2x Find the limit. Deltax->0 This is vague, because there's no "delta x" in the given f(x). Are you, by any chance, being asked to find the derivative of f(x)-(6x^2)-2x using the definition of the derivative?

OpenStudy (ralph):

Yes, i am.